If 2.8 kg of CO2 gas is stored in a 25 L tank at 20 °C, what is the required pressure of the tank?
Answer #1

here,

number of moles , n = 2.80 *10^3/44

n = 63.64 moles

temperature , T = 20 degree C

T = 293 K

Now , PV = n*R*T

P*25 = 63.64 * 0.0821 * 293

P = 61.2 atm

the required pressure of the tank is 61.2 atm